# Shared Spatial Memory Through Predictive Coding

### v1

@@ Methods
Emergence of hexagonal representations from geometric constraints. Finally, we justify why hexagonal activity patterns spontaneously emerge as the optimal solution. The path integration task fundamentally requires representing accumulated two-dimensional displacements in a compressed latent space with limited dimensionality. As established in Supplementary Information through Supplementary Proposition 1 and Theorem 1, any representation that faithfully performs path integration must satisfy translation equivariance: a physical displacement $\Delta\mathbf{r}$ must correspond to a consistent linear transformation in the latent space, independent of starting location. This geometric constraint, combined with the information efficiency requirement under the bottleneck architecture, necessitates a periodic encoding. Supplementary Theorem 2 proves that the minimal isotropic representation in two dimensions requires exactly three frequency directions uniformly distributed at $120^{\circ}$ angular separation, uniquely determined by first-order isotropy (zero mean) and second-order isotropy (identity-proportional metric). As demonstrated in Supplementary Corollary 3, these three directions generate effective six-fold symmetry due to cosine function parity: their superposition creates an interference pattern whose peak activations form a hexagonal lattice with $60^{\circ}$ rotational symmetry. Consequently, minimizing prediction error under equivariance and efficiency constraints causes LSTM hidden units to self-organize into grid-cell-like hexagonal patterns, providing the consistent spatial metric needed for path integration and forming the computational foundation for downstream prediction and planning.
@@
Having established a robust individual predictive model, the next challenge is to synchronize these models across agents. Central to this is a neural substrate designed to form a unified representation of the multi-agent system, which serves as the foundation for communication. We employ a dual-stream LSTM network that concurrently processes egocentric motion inputs (linear and angular velocities) from both the self-agent and its partner. The network is trained under a multi-faceted predictive objective: it must simultaneously predict its own future location, its partner’s future location, and, critically, the future Euclidean distance between them via a dedicated relational regression head. This compound predictive pressure compels the network’s shared latent representation to functionally specialize. This process gives rise to distinct neural populations, including units selectively tuned to the partner’s location—an artificial analogue of social place cells (SPCs; distinct from the self-position-encoding place cells described above, these units specifically encode partner locations within the observer’s reference frame), as detailed in our results (Fig. 4). This emergent social representation provides the rich, unified state, $S_{i,t}$, that informs the communication protocol. Furthermore, the explicit distance prediction serves as a critical gating mechanism for determining who and when communication is feasible, as detailed in Supplementary Method S1.4.1 (Social place cell module architecture in Table S11).
This architecture sets the stage for extending the predictive coding principle to the multi-agent domain, where we reformulate communication not as mere data transfer, but as a mechanism for collaborative prediction. The question shifts from “What information should I send?” to “What piece of my knowledge will maximally reduce my partner’s future uncertainty?” This is the essence of social predictive coding: agents learn to transmit only information that is novel and decision-relevant from the receiver’s perspective. To formalize this, we employ the information bottleneck (IB) principle. The goal is to learn a stochastic encoder, $p(m_{i,t}\mid S_{i,t})$, that maps the sender agent $i$’s state $S_{i,t}$ to a compressed message $m_{i,t}$. The optimal encoding must balance two competing objectives: maximizing the message’s predictive utility for the receiver while minimizing communication cost. Following the rate-distortion framework, we formulate this as minimizing a loss function that trades off distortion (predictive loss) against rate (compression cost):
@@
where $I(\cdot;\cdot)$ denotes mutual information, which is a measure of the mutual dependence between the two variables. $O_{j,t+1}$ is the future observation of the receiving agent $j$, $S_{j,t}$ is the receiver’s current state (serving as side information), and $\beta>0$ is a hyper-parameter balancing the trade-off. The Distortion term quantifies the negative of the information the message provides about the receiver’s future conditioned on what the receiver already knows—minimizing distortion thus maximizes predictive utility. The Rate term quantifies the information the message retains about the sender’s state—minimizing rate enforces compression. The parameter $\beta$ controls this trade-off: larger $\beta$ prioritizes compression, while smaller $\beta$ prioritizes predictive accuracy.
Directly optimizing Eq. (4) is intractable because computing mutual information requires integrating over high-dimensional and unknown data distributions. We therefore derive a tractable surrogate objective by constructing variational bounds for both the rate and distortion terms. We introduce a parametric encoder $q_{\varphi}(z|S_{i,t})$ that maps the sender’s state to a latent variable $z$ (representing the message $m=g(z)$, where $g(\cdot)$ is message decoder.), and a parametric decoder $p_{\vartheta}(O_{j,t+1}|z,S_{j,t})$ that predicts the receiver’s future observation.
First, we derive a tractable upper bound for the communication rate, $I(S_{i,t};m_{i,t})$. By introducing a fixed prior distribution $p(z)$ over the latent message space, the rate can be bounded by the Kullback-Leibler (KL) divergence between the approximate posterior and the prior:
@@
Since $H(O_{j,t+1}\mid S_{j,t})$ is constant, minimizing this bound is equivalent to minimizing the expected negative log-likelihood (reconstruction error). Combining these bounds yields the variational information bottleneck (VIB) objective, which is a tractable surrogate for the intractable IB loss. By substituting the upper bound for rate and the upper bound for distortion (omitting the constant $H(O_{j,t+1}\mid S_{j,t})$), we obtain:
@@
Each term in this loss directly implements a component of the social predictive coding framework. The first term (reconstruction loss) corresponds to distortion, minimizing prediction error and thereby maximizing the message’s utility for the receiver. The second term (KL divergence) corresponds to rate, enforcing compression and driving the emergence of efficient symbolic protocols. The hyperparameter $\beta$ enables systematic exploration of the rate-distortion trade-off, revealing how bandwidth constraints shape emergent communication structure (training configuration in Supplementary Table S10). The full derivation of the variational bounds is provided in Supplementary Method S1.6.
@@
The predictive coding framework culminates at the level of strategic decision-making. Having established models to predict its environment (Level 1) and its partners’ states (Level 2), the agent must now decide how to act in order to improve these predictive models over time. This transforms the agent from a passive observer into an active learner, a concept known as active inference. The optimal exploration strategy, from a predictive coding perspective, is to seek out experiences that maximally reduce the uncertainty of the agent’s internal generative model. In a large, partially-observable environment, this requires a principled approach to balance exploration and exploitation, a challenge we address with a hierarchical reinforcement learning (HRL) framework explicitly guided by predictive uncertainty.
This framework decomposes the complex navigation problem into two levels. The low-level controller is a deterministic planner (A* search) that executes concrete navigational sub-goals based on the agent’s current understanding of the shared world map. This offloads the complexities of pathfinding, allowing the high-level policy (meta-controller) to focus exclusively on the strategic question: “Where should I go next to learn the most?” The meta-controller is implemented as an actor-critic network and trained using multi-agent proximal policy optimization (MAPPO), a robust algorithm for cooperative settings. To fulfil this, each agent first partitions its occupancy grid into a $4\times 4$ regional summary, generating a 48-dimensional feature vector that includes the exploration ratio, walkability, and agent occupancy for each region. These features are fed into a shared actor-critic network with two fully connected layers (256 units, ReLU activation) to extract a common embedding. This embedding then branches into an actor head, which outputs a masked categorical distribution over the 16 regions for goal selection, and a critic head, which predicts scalar state values (Supplementary Table S12). The crucial insight is that the reward signal driving this high-level policy is not based on sparse, external rewards (like finding the target), but is instead generated intrinsically from the agent’s own state of uncertainty.
This is realized through an enhanced intrinsic curiosity module (ICM), which functions as the direct implementation of the active inference principle. Rather than relying on a learned forward dynamics model, our ICM directly estimates epistemic uncertainty by analyzing the geometry of the known-unknown boundary in the agent’s local map, augmented with social spatial information from the SPC module (Level 2). The composite intrinsic reward signal, $r_{t}^{\text{int}}$, is a weighted sum of three components, each reflecting a different facet of uncertainty reduction:
@@
Specifically, these components are implemented as follows: (1) The curiosity reward, $r_{\text{curiosity}}$, is a frontier-based score that directly rewards the selection of high-level goals in regions bordering unknown territory, where the shared predictive map is most uncertain. (2) The coordination reward, $r_{\text{coord}}$, promotes spatial division of labor by leveraging the inter-agent distance estimates $\hat{d}_{ij,t}$ provided by the SPC module. This component discourages redundant exploration by rewarding goal selections that maintain appropriate separation from teammates, directly operationalizing the principle that distributed exploration maximizes information gain. Critically, this establishes a computational dependency: the SPC module’s learned distance-tuned neurons (Fig. 4c) provide the essential geometric information that enables the ICM to generate coordination signals. (3) The exploration reward, $r_{\text{explore}}$, grants credit for discovering map cells that are unknown to the entire team, targeting states that are, by definition, maximally unpredictable. More details about ICM reward formulations and communication protocols are provided in Supplementary Method S1.4.
By training the MAPPO policy to maximize this rich, uncertainty-driven intrinsic reward, the HRL-ICM framework learns a sophisticated, coordinated exploration strategy from first principles. It does not rely on hand-crafted heuristics for exploration but instead learns to perform a form of Bayesian experimental design, constantly choosing actions to gather the most informative data. The entire system is trained end-to-end with the SPC module providing continuous geometric awareness of partner states, the ICM translating this into strategic exploration incentives, and the meta-controller selecting goals based on these incentives (hyperparameters and training configuration in Supplementary Table S13). Thus, from the lowest level of perception to the highest level of strategic planning, the entire architecture is unified under the overarching goal of building and refining a predictive model of the world by actively seeking out and resolving uncertainty.
@@
Z.F., Y.G. and J.W. conceived the predictive coding framework. Z.F., Y.G. Y.Z. and H.A. performed all experiments and data analysis. Y.G and H.A. contributed to the conception of the HRL-ICM framework and data collection. J.W. and Y.F. supervised the entire project. Y.W., and Y.F. wrote and reviewed the manuscript. All authors discussed the results and provided critical feedback on the manuscript. J.W. and Y.F. provided funding.
@@
S1.1 Grid cell network: Biological-inspired spatial representation
@@
The computational flow is designed to test a key principle of biological navigation: the anchoring of a dynamic, self-motion-based estimate of position to stable, absolute sensory cues. The process begins by with encoding the agent’s initial pose ($s_{0}$) into a high-dimensional, distributed representation using ensembles of virtual place and head direction (HD) cells [3]. This initial representation, grounded in an absolute reference frame, is projected through learnable linear layers to initialize the LSTM’s hidden state ($\mathbf{h}_{0}$) and cell state ($\mathbf{c}_{0}$). This initialization strategy is not a mere technical convenience; it is a critical design choice that simulates the neural mechanism of “remapping” or “resetting” observed when an animal enters a new environment. By grounding the integrator’s initial state in a veridical sensory observation, we prevent the unbounded accumulation of drift that is the primary failure mode of pure idiothetic integration.
@@
S1.1.2 Place and head direction cell ensembles as a sensory ground truth
@@
S1.1.3 Objective function for predictive spatial learning
@@
S1.1.4 Quantitative analysis of emergent spatial representations
@@
S1.2 Spatial memory generation module
The generation of a stable, allocentric spatial memory from a stream of egocentric visual inputs is a cornerstone of an agent’s individual world model. This process addresses a fundamentally ill-posed inverse problem: inferring a structured, top-down world representation from ambiguous, high-dimensional localized sensory data. Our architecture is designed as a hierarchical processing pipeline consisting of three primary stages: (1) object-centric feature extraction from the raw first-person view, (2) cross-view encoding of visual features into a latent representation, and (3) decoding into a structured bird’s-eye-view (BEV) spatial map [11, 10]. This pipeline leverages the complementary strengths of convolutional and transformer-based networks, and is trained end-to-end under a composite predictive objective that imposes a set of physical priors on the generative model, guiding it toward geometrically and photometrically plausible solutions. We describe each processing stage in detail below.
@@
S1.2.1 Target localization network
@@
S1.3 Emergent communication mechanism architecture
To operationalize the principle of Social Predictive Coding, we design a communication architecture that directly implements the variational information bottleneck (VIB) framework [1, 16]. The central hypothesis is that an efficient communication mechanism need not be manually designed but can emerge share when agents are optimized to transmit only the information that is maximally reductive of their partner’s future uncertainty. Consequently, the network architecture is not merely a tool for data compression, but a principled mechanism for learning a compact, structured, and task-relevant symbolic language from the ground up. This process is built upon a neural substrate capable of generating a rich, unified social representation, which serves as the input to the communication module. We first detail this substrate, followed by the communication architecture itself.
S1.3.1 Social representation substrate: Emergence of social place cells
Before an agent can decide what information to transmit, it must first form a comprehensive internal representation of the whole multi-agent system system [4]. This state, denoted as $S_{i,t}$ in our method, serves as the foundation for the entire communication mechanism. To this end, we design a specialized neural architecture, the Social Place Coding, to learn this rich social representation, as shown in the main text (Fig. 4a).
@@
S1.3.2 Encoder architecture: Implementing the VIB compression term
@@
S1.3.3 Decoder architecture: Implementing the VIB predictive utility term
@@
S1.3.4 VIB objective: Driving the emergence of efficient communication
@@
S1.4.1 Social place cell module for partner state estimation
The social place cell module serves as the perceptual foundation for multi-agent coordination by continuously estimating partner locations and computing inter-agent distances. This module extends the grid-cell-based path integration architecture described in Method S1.1 to the social domain through a dual-stream processing design.
@@
Both streams are initialized at episode start using the respective agents’ known starting poses, encoded through ensembles of virtual place cells and head-direction cells identical to those used in the individual path integration module (detailed in Method S1.1). The two LSTM hidden states, $\mathbf{h}_{\text{self},t}$ and $\mathbf{h}_{\text{partner},t}$, are then fused via element-wise summation to form a unified joint representation $\mathbf{h}_{\text{joint},t}=\mathbf{h}_{\text{self},t}+\mathbf{h}_{\text{partner},t}$. This shared representation is passed through a bottleneck layer (256 units, 50% dropout) that enforces a compressed, information-efficient encoding of the two-agent system state.

### v2

@@ Methods
Emergence of hexagonal representations from geometric constraints. Finally, we justify why hexagonal activity patterns spontaneously emerge as the optimal solution. The path integration task fundamentally requires representing accumulated two-dimensional displacements in a compressed latent space with limited dimensionality. As established in Supplementary Information through Supplementary Proposition 1 and Theorem 1, any representation that faithfully performs path integration must satisfy translation equivariance: a physical displacement $\Delta\mathbf{r}$ must correspond to a consistent linear transformation in the latent space, independent of starting location. This geometric constraint, combined with the information efficiency requirement under the bottleneck architecture, necessitates a periodic encoding. Supplementary Theorem 2 proves that the minimal isotropic representation in two dimensions requires exactly three frequency directions uniformly distributed at $120^{\circ}$ angular separation, uniquely determined by the first-order isotropy (zero mean) and second-order isotropy (identity-proportional metric). As demonstrated in Supplementary Corollary 3, these three directions generate effective six-fold symmetry due to cosine function parity: their superposition creates an interference pattern whose peak activations form a hexagonal lattice with $60^{\circ}$ rotational symmetry. Consequently, minimizing prediction error under equivariance and efficiency constraints causes LSTM hidden units to self-organize into grid-cell-like hexagonal patterns, providing the consistent spatial metric needed for path integration and forming the computational foundation for downstream prediction and planning.
@@
Having established a robust individual predictive model, the next challenge is to synchronize these models across agents. Central to this is a neural substrate designed to form a unified representation of the multi-agent system, which serves as the foundation for communication. We employ a dual-stream LSTM network that concurrently processes egocentric motion inputs (linear and angular velocities) from both the self-agent and its partner. The network is trained under a multi-faceted predictive objective: it must simultaneously predict its own future location, its partner’s future location, and, critically, the future Euclidean distance between them via a dedicated relational regression head. This compound predictive pressure compels the network’s shared latent representation to functionally specialize. This process gives rise to distinct neural populations, including units selectively tuned to the partner’s location—an artificial analogue of social place cells (SPCs; distinct from the self-position-encoding place cells described above, these units specifically encode partner locations within the observer’s reference frame), as detailed in our results (Fig. 4). This emergent social representation provides the rich, unified state, $S_{i,t}$, that informs the communication mechanism. Furthermore, the explicit distance prediction serves as a critical gating mechanism to determine who and when communication is feasible, as detailed in Supplementary Method S2.4.1 (Social place cell module architecture in Table S11).
This architecture sets the stage for extending the predictive coding principle to the multi-agent domain, where we reformulate communication not as mere data transfer, but as a mechanism for collaborative prediction. The question shifts from “What information should I send?” to “What piece of my knowledge will maximally reduce my partner’s future uncertainty?” This is the essence of social predictive coding: agents learn to share only information that is novel and decision-relevant from the receiver’s perspective. To formalize this, we employ the information bottleneck (IB) principle. The goal is to learn a stochastic encoder $p(m_{i,t}\mid S_{i,t})$ that maps sender agent $i$’s state $S_{i,t}$ to a compressed message $m_{i,t}$. The optimal encoding must balance two competing objectives: maximizing the message’s predictive utility for the receiver while minimizing communication cost. Following the rate-distortion framework, we formulate this as minimizing a loss function that balances distortion (predictive loss) and rate (compression cost):
@@
where $I(\cdot;\cdot)$ denotes mutual information, which is a measure of the mutual dependence between the two variables. $O_{j,t+1}$ is the future observation of the receiving agent $j$, $S_{j,t}$ is the receiver’s current state (serving as side information), and $\beta>0$ is a hyper-parameter balancing the trade-off. The Distortion term quantifies the negative of the information the message provides about the receiver’s future conditioned on what the receiver has already known—minimizing distortion thus maximizes predictive utility. The Rate term quantifies the information the message retains about the sender’s state—minimizing rate enforces compression. The parameter $\beta$ controls this trade-off: larger $\beta$ prioritizes compression, while smaller $\beta$ prioritizes predictive accuracy.
Directly optimizing Eq. (4) is intractable because computing mutual information requires integrating over high-dimensional and unknown data distributions. We therefore derive a tractable surrogate objective by constructing variational bounds for both the rate and distortion terms. We introduce a parametric encoder $q_{\varphi}(z|S_{i,t})$ that maps the sender’s state to a latent variable $z$ (representing the message $m=g(z)$, where $g(\cdot)$ is message decoder.), and a parametric decoder $p_{\vartheta}(O_{j,t+1}|z,S_{j,t})$ that predicts the receiver’s future observation. The full derivation of the variational bounds is provided in Supplementary Method S2.6.
First, we derive a tractable upper bound for the communication rate, $I(S_{i,t};m_{i,t})$. By introducing a fixed prior distribution $p(z)$ over the latent message space, the rate can be bounded by the Kullback-Leibler (KL) divergence between the approximate posterior and the prior according to Supplementary Lemma 4:
@@
Since $H(O_{j,t+1}\mid S_{j,t})$ is a constant, minimizing this bound is equivalent to minimizing the expected negative log-likelihood (reconstruction error). Combining these bounds yields the variational information bottleneck (VIB) objective, which is a tractable surrogate for the intractable IB loss. By substituting the upper bound for rate and the upper bound for distortion (omitting the constant $H(O_{j,t+1}\mid S_{j,t})$), we obtain a new loss function:
@@
Each term in this loss directly implements a component of the social predictive coding framework. The first term (reconstruction loss) corresponds to distortion, minimizing prediction error and thereby maximizing the message’s utility for the receiver. The second term (KL divergence) corresponds to rate, enforcing compression and driving the emergence of efficient symbolic mechanisms. The hyperparameter $\beta$ enables systematic exploration of the rate-distortion trade-off, revealing how bandwidth constraints shape emergent communication structure (training configuration in Supplementary Table S10).
@@
The predictive coding framework culminates at the level of strategic decision-making. Having established models to predict its environment (Level 1) and its partners’ states (Level 2), the agent must now decide how to act in order to improve these predictive models over time. This changes an agent from a passive observer into an active learner, a concept known as active inference. The optimal exploration strategy, from a predictive coding perspective, is to seek out experiences that maximally reduce the uncertainty of an agent’s internal generative model. In a large, partially-observable environment, this requires a principled approach to balance exploration and exploitation, a challenge we address with a hierarchical reinforcement learning (HRL) framework explicitly guided by predictive uncertainty.
This framework decomposes the complex navigation problem into two levels. The low-level controller is a deterministic planner (A* search) that executes concrete navigational sub-goals based on an agent’s current understanding of the shared world map. This offloads the complexities of pathfinding, allowing the high-level policy (meta-controller) to focus exclusively on the strategic question: “Where should I go next to learn the most?” The meta-controller is implemented as an actor-critic network and trained using multi-agent proximal policy optimization (MAPPO), a robust algorithm for cooperative settings. To fulfill this, each agent first partitions its occupancy grid into a $4\times 4$ regional summary, generating a 48-dimensional feature vector that includes the exploration ratio, walkability, and agent occupancy for each region. These features are fed into a shared actor-critic network with two fully connected layers (256 units with ReLU activation) to extract a common embedding. This embedding then branches into an actor head, which outputs a masked categorical distribution over the 16 regions for goal selection, and a critic head, which predicts scalar state values (Supplementary Table S12). The crucial insight is that the reward signal driving this high-level policy is not based on sparse, external rewards (like finding the target), but is instead generated intrinsically from an agent’s own state of uncertainty.
This is realized through an enhanced intrinsic curiosity module (ICM), which functions as the direct implementation of the active inference principle. Rather than relying on a learned forward dynamics model, our ICM directly estimates epistemic uncertainty by analyzing the geometry of the known-unknown boundary in an agent’s local map, augmented with social spatial information from the SPC module (Level 2). The composite intrinsic reward signal, $r_{t}^{\text{int}}$, is a weighted sum of three components, each reflecting a different facet of uncertainty reduction:
@@
Specifically, these components are implemented as follows: (1) The curiosity reward, $r_{\text{curiosity}}$, is a frontier-based score that directly rewards the selection of high-level goals in regions bordering unknown territory, where the shared predictive map is most uncertain. (2) The coordination reward, $r_{\text{coord}}$, promotes spatial division of labor by leveraging the inter-agent distance estimates $\hat{d}_{ij,t}$ provided by the SPC module. This component discourages redundant exploration by rewarding goal selections that maintain appropriate separation from teammates, directly operationalizing the principle that distributed exploration maximizes information gain. Critically, this establishes a computational dependency: the SPC module’s learned distance-tuned neurons (Fig. 4c) provide the essential geometric information that enables the ICM to generate coordination signals. (3) The exploration reward, $r_{\text{explore}}$, grants credit for discovering map cells that are unknown to the entire team, targeting states that are, by definition, maximally unpredictable. More details about ICM reward formulations and communication mechanisms are provided in Supplementary Method S2.4.
By training the MAPPO policy to maximize this rich, uncertainty-driven intrinsic reward, the HRL-ICM framework learns a sophisticated, coordinated exploration strategy from first principles. It does not rely on hand-crafted heuristics for exploration but instead learns to perform a form of Bayesian experimental design, constantly choosing actions to gather the most informative data. The entire system is trained end-to-end with the SPC module that provides continuous geometric awareness of partner states, the ICM that translates this into strategic exploration incentives, and the meta-controller that selects goals based on these incentives (hyperparameters and training configuration in Supplementary Table S13). Thus, from the lowest level of perception to the highest level of strategic planning, the entire architecture is unified under the overarching goal of building and refining a predictive model of the world by actively seeking out and resolving uncertainty.
@@
Z.F., Y.G. and J.W. conceived the predictive coding framework. Z.F., Y.G., Y.Z., and H.A. performed all experiments and data analysis. Y.G. and H.A. contributed to the conception of the HRL-ICM framework and data collection. J.W. and Y.F. supervised the entire project. Y.W. and Y.F. wrote and reviewed the manuscript. All authors discussed the results and provided critical feedback on the manuscript. J.W. and Y.F. provided funding.
@@
Appendix S1 System Overview
This section provides a comprehensive overview of our system architecture, training pipeline, and experimental configuration. The following detailed supplementary methods (Supplementary Method S2.1S2.4) describe individual components; here we explain how these components integrate into a unified framework for collaborative spatial navigation.
S1.1 System Architecture Overview
Our framework implements a modular architecture where specialized perceptual and cognitive components are composed into an end-to-end reinforcement learning system. The design philosophy follows a two-stage approach: first learning robust perceptual primitives independently, then integrating them for collaborative decision-making.
Pre-trained perceptual components.
Three core modules are trained independently before integration into the full multi-agent system:
  • Grid Cell Network (detailed in Supplementary Method S2.1): A recurrent neural network trained via self-supervised path integration to maintain a dynamic estimate of self-location from egocentric velocity commands. Through predictive learning alone, this module spontaneously develops hexagonally-periodic spatial representations analogous to biological grid cells, providing a stable metric foundation for navigation.
  • Spatial Memory Generation Module (detailed in Supplementary Method S2.2): A transformer-based encoder-decoder architecture that translates first-person RGB observations into structured bird’s-eye-view (BEV) spatial maps. Trained end-to-end on an egocentric-to-allocentric translation task, this module enables agents to construct persistent spatial memories from partial visual observations.
  • Emergent Communication Mechanism (detailed in Supplementary Method S2.3): A variational autoencoder (VAE) implementing the information bottleneck principle for bandwidth-constrained message passing. Trained to compress spatial maps while preserving predictive utility for partner agents, this module learns an efficient symbolic communication protocol without explicit supervision on message semantics.
Social representation substrate.
Building upon the individual grid cell network, a specialized Social Place Cell Module (detailed in Supplementary Method S2.4.1, Table S11) extends spatial coding to the multi-agent domain. This dual-stream LSTM architecture processes motion information from both self and partner agents, learning to predict not only individual positions but also inter-agent distance relationships. Through multi-task predictive training, this module develops functionally specialized neural populations—including units selectively tuned to partner locations (social place cells)—that provide the geometric awareness necessary for coordination.
End-to-end reinforcement learning framework.
The HRL-ICM Framework (detailed in Supplementary Method S2.4) integrates all pre-trained components into a hierarchical policy optimized for collaborative exploration and target search. A meta-controller network selects high-level exploration goals based on epistemic uncertainty, while an intrinsic curiosity module (ICM) generates dense reward signals from map coverage and inter-agent coordination. The social place cell module’s distance estimates enable the ICM to compute coordination rewards that encourage spatial division of labor, while the communication mechanism allows selective information sharing when bandwidth-efficient messages can reduce collective uncertainty. This entire system is trained end-to-end using Multi-Agent Proximal Policy Optimization (MAPPO) on the Memory-Maze benchmark.
The modular design enables interpretability (each component’s emergent representations can be analyzed independently), sample efficiency (pre-trained components provide strong initialization), and scalability (individual modules can be upgraded without full system retraining).
S1.2 Training Pipeline
The complete training procedure follows a three-stage curriculum designed to establish stable perceptual foundations before optimizing collaborative behavior:
  1. 1.
    Stage 1 – Individual Component Pre-training (Independent Training):
    • Grid Cell Network: Trained on synthetic trajectory datasets with randomized velocity sequences in $15\times 15$ meter environments. The network learns to predict place and head-direction cell activations after integrating motion commands over 100-step sequences. Training uses Adam optimizer (learning rate $1\times 10^{-3}$, batch size 64) for approximately 50 epochs until gridness scores stabilize (typically >0.3>0.3).
    • BEV Prediction Network: Trained on paired egocentric RGB / allocentric BEV map datasets collected from agent trajectories in Memory-Maze environments. The transformer encoder-decoder is optimized to minimize a composite loss balancing occupancy prediction (binary cross-entropy), appearance reconstruction (masked MSE), and spatial smoothness. Training uses Adam optimizer (learning rate $1\times 10^{-4}$, batch size 64) until validation IoU exceeds 0.85.
    • Communication VAE: Trained on spatial map reconstruction tasks with information bottleneck objective. Paired sender/receiver map samples are generated from dual-agent exploration trajectories. The encoder-decoder architecture is optimized to balance reconstruction accuracy against KL divergence from a Gaussian prior ($\beta=1.0$). Training uses Adam optimizer (learning rate $1\times 10^{-3}$, batch size 32) for 50 epochs, achieving 32:1 compression with $5\%$ reconstruction error.
  2. 2.
    Stage 2 – Social Place Cell Module Training (Dual-Agent Path Integration):
    • The social place cell module is initialized with pre-trained grid cell network weights for the self-stream, then extended with a partner-stream LSTM.
    • Training data consists of synchronized dual-agent trajectories where both agents’ poses and velocities are recorded. The partner velocity is estimated from visual observations of the partner’s motion when visible.
    • The module is trained to predict final-timestep place/head-direction activations for both agents plus their Euclidean distance via multi-task loss: $\mathcal{L}_{\text{SPC}}=\mathcal{L}_{\text{self}}+\mathcal{L}_{\text{partner}}+\mathcal{L}_{\text{distance}}$.
    • Training uses Adam optimizer (learning rate $1\times 10^{-3}$, batch size 64) for 30 epochs. Success criterion: distance prediction MAE <1.0<1.0 grid cells and emergence of distinct social place cell populations (verified via neural selectivity analysis).
  3. 3.
    Stage 3 – HRL-ICM End-to-End Training (Multi-Agent Reinforcement Learning):
    • Pre-trained components (grid cell, BEV prediction, communication VAE, social place cell module) are integrated into the full HRL-ICM system.
    • Initial Freezing Phase (first $10^{4}$ environment steps): All pre-trained components have frozen parameters to allow meta-controller and ICM to stabilize against fixed perceptual representations. Only the meta-controller actor-critic network and communication gating policy are updated via MAPPO.
    • Joint Fine-tuning Phase (remaining training): All components are unfrozen and jointly optimized end-to-end. Meta-controller updates continue via MAPPO (learning rate $3\times 10^{-4}$), while perceptual modules receive gradients from both task rewards and their original supervised objectives (weighted $0.1\times$ to prevent catastrophic forgetting).
    • Training environments: Procedurally generated Memory-Maze instances with increasing complexity (maze sizes $15\times 15$ to $39\times 39$, 2–5 agents). Training continues for $5\times 10^{6}$ environment steps or until success rate on $29\times 29$ mazes exceeds 70%.
    • Curriculum: Maze size increases every $10^{6}$ steps; communication bandwidth decreases from 128 bits/step to target bandwidth (4–16 bits/step) over first $2\times 10^{6}$ steps to encourage robust communication strategies.
This staged training strategy ensures that higher-level reasoning builds upon stable low-level representations, preventing the RL policy from corrupting perceptual modules during exploration. The modular architecture also enables ablation studies where individual components can be selectively frozen or replaced.
S1.3 Experimental Configuration
The complete experimental setup encompasses hardware infrastructure, simulation environment parameters, and training hyperparameters across all system components.
Evaluation protocol.
All reported results are averaged over 100 evaluation episodes with fixed random seeds for reproducibility. For bandwidth scaling experiments (Fig. 3c–d), agents are tested across latent dimensions $z_{\text{dim}}\in\{16,32,64,128,256\}$ corresponding to $\{4,8,16,32,64\}$ bits/step (assuming 4-bit quantization). For generalization tests, models trained on $29\times 29$ mazes are evaluated on unseen $39\times 39$ mazes without further training. Statistical significance is assessed via two-sample $t$-tests with Bonferroni correction for multiple comparisons.
Forward reference.
Detailed specifications for each component’s network architecture, loss functions, and training procedures are provided in the following supplementary methods sections: Grid Cell Network (Supplementary Method S2.1, Tables S1, S2), Spatial Memory Generation (Supplementary Method S2.2, Tables S3, S4, S5), Communication Mechanism (Supplementary Method S2.3, Tables S8, S9, S10), and HRL-ICM Framework (Supplementary Method S2.4, Tables S11, S12, S13).
@@
S2.1 Grid cell network: Biological-inspired spatial representation
@@
The computational flow is designed to test a key principle of biological navigation: the anchoring of a dynamic, self-motion-based estimate of position to stable, absolute sensory cues. The process begins by encoding the agent’s initial pose ($s_{0}$) into a high-dimensional, distributed representation using ensembles of virtual place and head direction (HD) cells [3]. This initial representation, grounded in an absolute reference frame, is projected through learnable linear layers to initialize the LSTM’s hidden state ($\mathbf{h}_{0}$) and cell state ($\mathbf{c}_{0}$). This initialization strategy is not a mere technical convenience; it is a critical design choice that simulates the neural mechanism of “remapping” or “resetting” observed when an animal enters a new environment. By grounding the integrator’s initial state in a veridical sensory observation, we prevent the unbounded accumulation of drift that is the primary failure mode of pure idiothetic integration.
@@
S2.1.2 Place and head direction cell ensembles as a sensory ground truth
@@
S2.1.3 Objective function for predictive spatial learning
@@
S2.1.4 Quantitative analysis of emergent spatial representations
@@
S2.2 Spatial memory generation module
The generation of a stable, allocentric spatial memory from a stream of egocentric visual inputs is a cornerstone of an agent’s individual world model. This process addresses a fundamentally ill-posed inverse problem: inferring a structured, top-down world representation from ambiguous, high-dimensional localized sensory data. Our architecture is designed as a hierarchical processing pipeline consisting of three primary stages: (1) object-centric feature extraction from the raw first-person view; (2) cross-view encoding of visual features into a latent representation; and (3) decoding into a structured bird’s-eye-view (BEV) spatial map [11, 10]. This pipeline leverages the complementary strengths of convolutional and transformer-based networks, and is trained end-to-end under a composite predictive objective that imposes a set of physical priors on the generative model, guiding it toward geometrically and photometrically plausible solutions. We describe each processing stage in detail below.
@@
S2.2.1 Target localization network
@@
S2.3 Emergent communication mechanism architecture
To operationalize the principle of Social Predictive Coding, we design a communication architecture that directly implements the variational information bottleneck (VIB) framework [1, 16]. The central hypothesis is that an efficient communication mechanism need not be manually designed but can emerge when agents are optimized to transmit only the information that is maximally reductive of their partner’s future uncertainty. Consequently, the network architecture is not merely a tool for data compression, but a principled mechanism for learning a compact, structured, and task-relevant symbolic language from the ground up. This process is built upon a neural substrate capable of generating a rich, unified social representation, which serves as the input to the communication module. We first detail this substrate, followed by the communication architecture itself.
S2.3.1 Social representation substrate: Emergence of social place cells
Before an agent can decide what information to transmit, it must first form a comprehensive internal representation of the whole multi-agent system [4]. This state, denoted as $S_{i,t}$ in our method, serves as the foundation for the entire communication mechanism. To this end, we design a specialized neural architecture, the Social Place Coding, to learn this rich social representation, as shown in the main text (Fig. 4a).
@@
S2.3.2 Encoder architecture: Implementing the VIB compression term
@@
S2.3.3 Decoder architecture: Implementing the VIB predictive utility term
@@
S2.3.4 VIB objective: Driving the emergence of efficient communication
@@
S2.4.1 Social place cell module for partner state estimation
The social place cell module serves as the perceptual foundation for multi-agent coordination by continuously estimating partner locations and computing inter-agent distances. This module extends the grid-cell-based path integration architecture described in Method S2.1 to the social domain through a dual-stream processing design.
@@
Both streams are initialized at episode start using the respective agents’ known starting poses, encoded through ensembles of virtual place cells and head-direction cells identical to those used in the individual path integration module (detailed in Method S2.1). The two LSTM hidden states, $\mathbf{h}_{\text{self},t}$ and $\mathbf{h}_{\text{partner},t}$, are then fused via element-wise summation to form a unified joint representation $\mathbf{h}_{\text{joint},t}=\mathbf{h}_{\text{self},t}+\mathbf{h}_{\text{partner},t}$. This shared representation is passed through a bottleneck layer (256 units, 50% dropout) that enforces a compressed, information-efficient encoding of the two-agent system state.